One number is four times more than another. The product of the numbers is 96. Find the numbers.
Please help! Thank you!

Answer by LinnW(1048)

x = first number
y = second number
x = y+4
x*y=96
substitute (4+y) for x in above
(4+y)*y = 96
4y + y^2 = 96
subtract 96 from each side
y^2 + 4y -96 = 0
Since 8 x 12 = 96 we can factor
(y+12)(y-8)=0
So y = -12 or y = 8
since x = y+4
for y = -12 , x = -8
for y = 8 , x = 12
The possible values for the two numbers are 8,12 or -8,-12